OpenStudy (anonymous):

Which of the following options is an equivalent function to f(x) = 3(2)3x?

OpenStudy (anonymous):

Answers: f(x) = 3(8)x f(x) = 24x f(x) = 27(8)x f(x) = 3(8x)

OpenStudy (anonymous):

@Vocaloid

OpenStudy (danjs):

none

OpenStudy (anonymous):

@DanJS what do u mean

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

@Saiyan @CBARREDO1 @Vocaloid please help

OpenStudy (anonymous):

@Saiyan

OpenStudy (danjs):

As typed, you have f(x) = 3 * 2 * 3*x = 18x no way to get any of those answers with that...

OpenStudy (anonymous):

@DanJS sry its 3(2)^3x

OpenStudy (cbarredo1):

f(x)= 3(8x)

OpenStudy (anonymous):



OpenStudy (cbarredo1):

The answer is f(x)= 3(8x)

OpenStudy (anonymous):

@DanJS

OpenStudy (cbarredo1):

THE ANSWER IS: f(x)= 3(8x)

OpenStudy (anonymous):

to the power of 3x

OpenStudy (anonymous):

3(2)^3x

OpenStudy (cbarredo1):

@jaredhm29 THE ANSWER IS: f(x)= 3(8x)

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

it is A \[ 3(8)^x\]

OpenStudy (cbarredo1):

What have I been saying?!

OpenStudy (anonymous):

Which of the following options results in a graph that shows exponential decay? f(x) = 0.6(2)x f(x) = 3(0.7)x f(x) = 0.4(1.6)x f(x) = 20(3)x

OpenStudy (mathstudent55):

\(f(x) = 3(2)^{3x}\) Rule: To raise a power to a power, multiply powers. \(\large (a^m)^n = a^{mn}\) \(\large f(x) = 3(2)^{3x}\) Apply the rule above in reverse to \(2^{3x} \) \(\large f(x) = 3(2^3)^x\) \(\large f(x) = 3 \cdot 8^x \)

OpenStudy (anonymous):

Can u help Which of the following options results in a graph that shows exponential decay? f(x) = 0.6(2)x f(x) = 3(0.7)x f(x) = 0.4(1.6)x f(x) = 20(3)x @mathstudent55

OpenStudy (cbarredo1):

f(x) = 3(0.7)x

OpenStudy (anonymous):

Which of the following options results in a graph that shows exponential growth? f(x) = 0.4(3)x f(x) = 3(0.5)x f(x) = 0.8(0.9)x f(x) = 0.9(5)−x

OpenStudy (cbarredo1):

f(x) = 0.4(3)x

OpenStudy (anonymous):

For f(x) = 0.01(2)x, find the average rate of change from x = 3 to x = 8. 0.08 0.496 2.48 5

OpenStudy (anonymous):

@CBARREDO1

OpenStudy (mathstudent55):

\(\large f(x) = ak^x\) As exponents become larger, If \(k \gt 1\), it is exponential growth. If \(0 \lt k \lt 1\) it is exponential decay.

OpenStudy (anonymous):

oh thx

OpenStudy (mathstudent55):

Find f(8) and f(3). Then do (f(8) - f(3)) / (8 - 3)

OpenStudy (anonymous):

can u help me with these 2 For f(x) = 0.01(2)x, find the average rate of change from x = 3 to x = 8. 0.08 0.496 2.48 5 and A wrestling match has eliminations each round. The table below shows the number of wrestlers in each of the first 4 rounds of the tournament: Round (x) 1 2 3 4 Wrestlers f(x) 64 32 18 9 Compute the average rate of change of f(x) from x = 1 to x = 4 and identify the meaning of that rate. 55; there were 55 fewer wrestlers between rounds 1 and 4 3; there were 3 rounds between rounds 1 and 4 negative fifty five over three; on average, there were 55 fewer players for every 3 rounds negative three over fifty five; on average, there were 3 fewer players for every 55 rounds

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@CBARREDO1

OpenStudy (mathstudent55):

Round x 1 4 Wrestlers f(x) 64 9 f(1) = 64 f(4) = 9 Subtract f(4) - f(1) and divided by the difference (4 - 1)

OpenStudy (anonymous):

So B?

OpenStudy (anonymous):

@mathstudent55 what about number 1